Find the perimeter of an isosceles triangle whose base is 40 cm and whose base angle is 70o


1
Expert's answer
2021-11-23T17:58:52-0500

since the triangle is isosceles, the other two sides will have equal length. Let call the length of each side "x" .

Draw a perpendicular from vertex to base

(Bisector of base)



You will get a right triangle with a leg "20 cm" and one angle "70 degrees"


Hence can use "cosine" "ratio" to obtain other side.


"cos(70°)=\frac{20}{x}"


"x=\frac{20}{cos(70°)}=58.48cm"


Add up the sides to get perimeter


Perimeter "= 58.48+58.48+40=156.96"